OpenStudy (mackenzie2013):

Tia's teacher asked her to find the product of 8 and 207 in her head. Which of the following describes the best way for Tia to mentally calculate the answer? Multiply 8 by 200 and subtract 8 times 7. Multiply 8 by 200 and add 8 times 7. Multiply 8 by 210 and subtract 8 times 7.

hartnn (hartnn):

207*8 = (200+7)*8 = 200*8 +7*8

OpenStudy (desiraem):

multiply 8 by 200 and add 8 times 7.